Statement: “The product of an even number and any other number is even.”
Using direct proof method.
1
Expert's answer
2020-11-08T18:37:41-0500

Suppose n is even integer. Then by the definition of even numbers, n = 2k for some integer k.

Suppose m is an integer.

Then by substitution we have "m\cdot n=m(2k)=2(mk)=2q" for some integer "q=mk". Therefore by the definition of even numbers the product "m\cdot n" is an even number.

 This completes the proof.

Therefore, the product of an even number and any other number is even.